# Defending Against Malicious Behaviors in Federated Learning with Blockchain

### v1

@@ 4.5 Theoretical Analysis on Malicious Votes
When all clients are rational and there is no collusion among malicious clients, a malicious client should not make a malicious vote.
@@
Under A4, each client makes an independent decision on voting. Let $r$ be the ratio of malicious clients in $\mathcal{K}_{v}$, there are $r\cdot n_{v}$ malicious clients in $\mathcal{K}_{v}$ and $(1-r)\cdot n_{v}$ honest clients. If $r\cdot n_{v}<(1-r)\cdot n_{v}$, i.e. missing $r<0.5$, each malicious client will lose $\gamma_{v}$; if $r\cdot n_{v}>(1-r)\cdot n_{v}$, i.e. missing $r>0.5$, each malicious client will gain $\frac{(1-r)\cdot n_{v}\cdot\gamma_{v}}{r\cdot n_{v}}=\frac{1-r}{r}\gamma_{v}$. The expected return $\mathcal{R}$ of a malicious client will be
@@
Additionally, there is a game among malicious clients. As Eq. (4) is known by all clients in advance. A malicious client can easily make an honest vote to gain tokens from other malicious clients who make dishonest votes. However, according to Nash Equilibrium [41], we are certain that, in the long run, no malicious clients will make dishonest votes.
@@
In practice, A4 can be relaxed, where multiple malicious clients work together to attach the FL system. If A2 holds, we are certain that the malicious voters will reach a consensus internally before they act to win the majority vote. Intuitively, all malicious voters can be considered as a group together. In this case, this ”group” will behave exactly as the single malicious client in Theorem 1 based on the same reasoning. The proof is omitted.
@@ 5 Experiments
(a) $\eta$ = 0.1.
(b) $\eta$ = 0.2.
(c) $\eta$ = 0.3.
(d) $\eta$ = 0.4.
Figure 5: Federated training under different values of the ratio of malicious clients ($\eta$). The solid lines are the mean AUROCs and the shaded regions are 1 standard deviation around the means. FedAVG w/ Block significantly outperforms FedAVG w/ mal, while being comparable with FedAVG w/o mal, the performance upper bound under this setup.
(a) $\eta$ = 0.1.
(b) $\eta$ = 0.2.
(c) $\eta$ = 0.3.
(d) $\eta$ = 0.4.
@@
5.1 Experimental Setup
We evaluate the proposed framework in a simulated environment.
5.1.1 Data and Task
We consider a standard classification task, namely loan default prediction. We use the Kaggle Lending Club dataset [^4] https://www.kaggle.com/datasets/wordsforthewise/lending-club to simulate a realistic financial application scenario. We pre-process the raw dataset by dropping all entries with missing values. For the labels, we only keep “Fully Paid” and “Charged Off” to simplify the task as a binary classification task. We randomly select $80\%$ of the data as the training set and use the rest of the data as the test set. The training set is split into $K$ subsets of equal size and distributed across $K$ clients. Within each client, $20\%$ of the local data are randomly selected as the validation set.
5.1.2 Implementation
There are $K=50$ clients in the system and each client is initialized with 64 tokens. We use a 3-layer multi-layer perceptron (MLP) as the network backbone. Apart from the last layer, each layer of the MLP has 128 hidden nodes. We use a standard Adam [42] optimizer with fixed learning rate $10^{-3}$ and batch size 128. No data augmentation is applied. We use the binary accuracy as both the local validation score and evaluation metric. We consider a simple data poisoning attack [14], where malicious clients are trained to confuse the model. All baselines are implemented in PyTorch 1.12.1 [43] on one NVIDIA Tesla T4 GPU. We leverage Ethereum smart contracts to deploy our reward-and-slash design in a private blockchain and simulate the training process using the Python library Web3.py [^5] https://web3py.readthedocs.io/en/v5/. We set $\epsilon=0.05$ based on empirical experience. [^6] We notice that too small $\epsilon$ can cause large oscillation, which slows the convergence, and too large $\epsilon$ can facilitate the convergence at the expense of decreased detection performance, i.e. missing the system fails to remove the majority of malicious clients.
5.1.3 Baselines
We consider 4 baselines. The first one is an Oracle approach, a centralized baseline without malicious attacks. The Oracle should provide the upper-bound performance of the experiment. The second one is FedAVG without malicious attacks (denoted as FedAVG w/o mal), which is equivalent to FedAVG under $\eta=0$ and should provide the upper-bound performance for a decentralized environment. The third one is FedAVG under malicious attacks (denoted as FedAVG w/ mal), where $\eta$ of clients are malicious. The fourth one is the proposed method, FedAVG with blockchain under malicious attacks (denoted as FedAVG w/ block). For FL baselines, $10\%$ of clients are randomly selected to perform local training at each epoch. For FedAVG w/ block, we simply use the remaining $90\%$ of the clients as voters.
@@ 5.2.1 Empirical Analysis on Malicious Voters
To empirically validate the theoretical result in Sec. 4.5, we first simulate a hypothetical scenario where there are only honest proposers. As there are more honest proposers than malicious proposers at each round on average, the effect of malicious weights can be seen as slowing the convergence and decreasing the global performance, which will be validated in Sec. 5.1.3. Note, due to A4, we further simplify the scenario to focus on the behavior of malicious voters. As shown in Fig. 4, given the set of the hyperparameter for slashing voters $\gamma_{r}=\{2,4,8,16,32\}$, the malicious voters will be eliminated from the system shortly (i.e. missing their average tokens decline to 00 within $\approx 40$ epochs).
@@ 5.2.3 Analysis of Token Distributions
Fig. 7, 8 and 9 depict the average tokens remaining in honest and malicious proposers during the FL training process. We observe that honest proposers gradually accumulate more tokens while malicious proposers have fewer tokens over sufficient training epochs. Eventually, most malicious proposers lose the ability to participate in staking and are removed from the FL system, as their remaining tokens are insufficient. This meets the expectations of our system design.
@@ 5.2.6 Limitations
In this work, as the experimental results aim to evaluate the robustness of the proposed framework, several practical challenges are simplified, e.g. missing staleness [44], storage, and privacy [18]. Further, the proposed method requires more computational power than traditional methods due to mining (blockchain computing) and voting. Finally, large models have gained in popularity in practical applications, e.g. missing ViT [45] and GPT-3 [46]. This raises the question on how to efficiently handle on-chain aggregation for large models. Future work thus will aim to address these limitations to facilitate the research and development of FL with blockchain.
@@ 6 Conclusion
In this work, we explore an under-explored research direction, namely using FL and blockchain to defend against malicious behaviors. The defense mechanism is twofold. We use on-chain smart contracts to replace the traditional central server and propose a stake-based majority voting mechanism to detect client-side malicious behaviors. We not only provide a solution to the problem of interest, an emerging direction on trustworthy ML, but also show the robustness of the proposed method and provide the first empirical understanding of the problem.
@@ Acknowledgment
The authors would like to thank Shuhao Zheng from the School of Computer Science, McGill University for the discussion in the early stage.

### v2

@@ 4.5 Theoretical Analysis on Malicious Votes
When all clients are rational, a malicious client should not make a malicious vote.
@@
Let us consider a multi-agent scenario, where malicious clients can collude. No matter how the malicious clients cooperate, there are two types of malicious clients. The first type behaves maliciously to achieve the goal of sabotaging the FL training, i.e. missing, lowering the global performance. The second type acts honestly to hide themselves and survive to be able to implement the complex policy to sabotage the FL training at the a later stage. If the malicious clients belong to the second type, they are factually “honest” ones.
Let $r$ be the ratio of malicious clients in $\mathcal{K}_{v}$, there are $r\cdot n_{v}$ malicious clients in $\mathcal{K}_{v}$ and $(1-r)\cdot n_{v}$ honest clients. If $r\cdot n_{v}<(1-r)\cdot n_{v}$, i.e. missing $r<0.5$, each malicious client will lose $\gamma_{v}$; if $r\cdot n_{v}>(1-r)\cdot n_{v}$, i.e. missing $r>0.5$, each malicious client will gain $\frac{(1-r)\cdot n_{v}\cdot\gamma_{v}}{r\cdot n_{v}}=\frac{1-r}{r}\gamma_{v}$. The expected return $\mathcal{R}$ of a malicious client will be
@@
If A2 holds, we are certain that the malicious voters will reach a consensus internally before they act to win the majority vote. Intuitively, all malicious voters can be considered as a group together. In this case, this ”group” will behave exactly as the single malicious client in Theorem 1 based on the same reasoning. The proof is omitted.
@@ 4.6 Training
4.7 Computation Complexity and Cost
For proposers, the primary computational cost is driven by the local training algorithm, mirroring the structure of traditional FL. For voters, their computation cost stems from evaluating the aggregated global model. In addition to mining blocks, miners engage in further computations by consolidating global model updates, akin to the responsibilities of a centralized aggregator in traditional FL. The overall computational complexity is contingent upon the underlying training network backbone.
The communication and storage costs are outlined as follows. We assume the model size is $M$. During an epoch, the communication cost for a client (i.e. missing, a proposer or a voter) is $O(M)$. And the storage cost on the blockchain is $O(K\cdot M)$, where $K$ is the number of clients.
@@ 5 Experiments
Figure 5: Federated training under different values of the ratio of malicious clients ($\eta$). Each subfigure shows the training AUROCs when the training time (i.e. missing the number of epochs) increases. The solid lines are the mean AUROCs and the shaded regions are 1 standard deviation around the means. We compare the performances of FedAVG with blockchain (i.e., w/ Block), FedAVG with malicious clients (i.e., w/ mal), and FedAVG without malicious clients (i.e., w/o mal). We observe that FedAVG w/ Block significantly outperforms FedAVG w/ mal, while being comparable with FedAVG w/o mal, the performance upper bound under this setup.
5.1 Experimental Setup
We first evaluate the proposed framework in a simulated environment.
5.1.1 Data and Task
We consider a standard binary classification task, namely loan default prediction. We use the Kaggle Lending Club dataset [^4] https://www.kaggle.com/datasets/wordsforthewise/lending-club to simulate a realistic financial application scenario. We pre-process the raw dataset by dropping all entries with missing values. For the labels, we only keep “Fully Paid” and “Charged Off” to simplify the task as a binary classification task. We randomly select $80\%$ of the data as the training set and use the rest of the data as the test set. The training set is split into $K$ subsets of equal size and distributed across $K$ clients. Within each client, $20\%$ of the local data are randomly selected as the validation set.
5.1.2 Implementation
There are $K=50$ clients in the system and each client is initialized with 64 tokens. We use a 3-layer multi-layer perceptron (MLP) as the network backbone. Apart from the last layer, each layer of the MLP has 128 hidden nodes. We use a standard Adam [46] optimizer with fixed learning rate $10^{-3}$ and batch size 128. No data augmentation is applied. We use the binary accuracy as both the local validation score and evaluation metric. In our experiments, malicious clients are randomly selected before the training according to the ratio $\eta$. In each training round, if a malicious client is selected as the proposer or voter, it will act maliciously as described in Sec. 4.555If a malicious client acts honestly, then it will be considered as an honest one and makes no harm to the system. We consider a simple data poisoning attack [14], where malicious clients are trained to confuse the model. Specifically, the local models are trained to lower the model performance by using the wrong labels, but maintaining a low weight divergence from the aggregated weights from the last round. All baselines are implemented in PyTorch 1.12.1 [47] on one NVIDIA Tesla T4 GPU. We leverage Ethereum smart contracts to deploy our reward-and-slash design in a private blockchain and simulate the training process using the Python library Web3.py [^6] https://web3py.readthedocs.io/en/v5/. We set $\epsilon=0.05$ based on empirical experience. [^7] We notice that too small $\epsilon$ can cause large oscillation, which slows the convergence, and too large $\epsilon$ can facilitate the convergence at the expense of decreased detection performance, i.e. missing the system fails to remove the majority of malicious clients.
5.1.3 Baselines
So far, there is no such blockchain-based FL baseline suitable for the comparative evaluation of counteracting malicious behaviors in our problem setting. It is worth mentioning that the proposed framework can be integrated with the existing FL method. In our experiments, we use FedAVG as the backbone FL method to illustrate our defending mechanism. We consider 4 baselines. The first one is an Oracle approach, a centralized baseline without malicious attacks. The Oracle should provide the upper-bound performance of the experiment. The second one is FedAVG without malicious attacks (denoted as FedAVG w/o mal), which is equivalent to FedAVG under $\eta=0$ and should provide the upper-bound performance for a decentralized environment. The third one is FedAVG under malicious attacks (denoted as FedAVG w/ mal), where $\eta$ of clients are malicious. The fourth one is the proposed method, FedAVG with blockchain under malicious attacks (denoted as FedAVG w/ block). For FL baselines, $10\%$ of clients are randomly selected to perform local training at each epoch. For FedAVG w/ block, we simply use the remaining $90\%$ of the clients as voters.
(a) $\eta$ = 0.1.
(b) $\eta$ = 0.2.
(c) $\eta$ = 0.3.
(d) $\eta$ = 0.4.
@@ 5.1.3 Baselines
(a) $\eta$ = 0.1.
(b) $\eta$ = 0.4.
(c) $\eta$ = 0.1.
(d) $\eta$ = 0.4.
Figure 7: (a)-(b) Number of award and slash epochs for clients when setting the parameter for slashing proposers as $\gamma_{p}=8$. (c)-(d) The number of slash epochs increases when the number of malicious proposers increases.
@@ 5.2.1 Empirical Analysis on Malicious Voters
To empirically validate the theoretical result in Sec. 4.5, we first simulate a hypothetical scenario where there are only honest proposers. As there are more honest proposers than malicious proposers at each round on average, the effect of malicious weights can be seen as slowing the convergence and decreasing the global performance, which will be validated in Sec. 5.1.3. Here, we further simplify the scenario to focus on the behavior of malicious voters. As shown in Fig. 4, given the set of the hyperparameter for slashing voters $\gamma_{r}=\{2,4,8,16,32\}$, the malicious voters will be eliminated from the system shortly (i.e. missing their average tokens decline to 00 within $\approx 40$ epochs).
@@ 5.2.3 Analysis of Token Distributions
Fig. 6 depicts the average tokens remaining in honest and malicious proposers during the FL training process when $\gamma_{p}=8$. We observe that honest proposers gradually accumulate more tokens while malicious proposers own fewer tokens as training progresses. Eventually, most malicious proposers lose the eligibility to participate in staking and are removed from the FL system, as their remaining tokens are insufficient. Looking more closely at the case where few malicious proposers are in the system (Fig. 6(a)), we note that proposals are initially accepted until a revert and slash step is performed around epoch 32 (indicated by the sudden drop in tokens for malicious proposers). As more malicious proposers are in the system (Fig. 6(d)), revert and slash steps occur more frequently. This is further highlighted by Fig. 7(a) and 7(b), which demonstrates for each epoch if it corresponds to an award or a revert and slash step. Fig. 7(c) and 7(d) further provide the cumulative sum over the number of award and slash epochs, illustrating that the majority of the epochs consist of award epochs and that the fraction of revert and slash episodes increases with the rising fraction of malicious proposers. Finally, we depict the average tokens remaining in honest and malicious proposers under various configurations of $\gamma_{p}$ in Fig. 8 and Fig. 9. We observe the same phenomenon as Fig. 6, which aligns with the expectations of our system design and reinforces the effectiveness of our approach.
@@ 5.2.6 Limitations
In this work, as the experimental results aim to evaluate the robustness of the proposed framework, several practical challenges are simplified, e.g. missing staleness [48], storage, and privacy [44]. Further, the proposed method requires more computational power than traditional methods due to mining (blockchain computing) and voting. Finally, large models have gained in popularity in practical applications, e.g. missing ViT [49] and GPT-3 [50]. This raises the question of how to efficiently handle on-chain aggregation for large models. Future work thus will aim to address these limitations to facilitate the research and development of FL with blockchain.
5.3 Cost Analysis
In our experiments, we adopt the Kaggle Lending Club dataset [^8] https://www.kaggle.com/datasets/wordsforthewise/lending-club to simulate a realistic financial application scenario. The total client number is $K=50$. For each client, the generated model update file is with a size of $587$KB. Therefore, the communication cost incurred by a proposer or a voter is $587$KB and the storage cost on the blockchain is $587\times 50$KB $=28.66$MB.
6 Additional Experiments
In Sec. 5, we examine the performance of the proposed system on a simple learning task on IID data. In this section, we further evaluate the robustness of the proposed system with a more complex task on non-IID data.
(a) $\eta=0.1$
(b) $\eta=0.2$
(c) $\eta=0.3$
(d) $\eta=0.4$
Figure 13: Federated training under different values of ratio of malicious clients ($\eta$). The solid lines are the mean AUROCs and the shaded regions are 1 standard deviation around the means.
Table 4: Performance comparison under different values of ratio of malicious clients ($\eta$). The reported numbers are mean and standard deviation under 5 random seeds.
Model $\eta=0.1$ $\eta=0.2$ $\eta=0.3$ $\eta=0.4$
FA w/ M $0.677\pm 0.014$ $0.640\pm 0.012$ $0.594\pm 0.030$ $0.483\pm 0.080$
FA w/ B $0.683\pm 0.008$ $0.688\pm 0.024$ $0.676\pm 0.023$ $0.673\pm 0.021$
FA w/o M $0.695\pm 0.004$ $0.695\pm 0.004$ $0.695\pm 0.004$ $0.695\pm 0.004$
Oracle $0.733\pm 0.014$ $0.733\pm 0.014$ $0.733\pm 0.014$ $0.733\pm 0.014$
6.1 Experimental Setup
6.1.1 Data and Task
We consider a standard multi-label classification (MLC) task [51] to simulate a realistic clinical application scenario. We use the ChestX-ray14999https://nihcc.app.box.com/v/ChestXray-NIHCC dataset [20] and leverage the first $6{\times}10^{4}$ chest X-ray images (CXRs) as our non-IID dataset. We use $80\%$ of the data as the training set and use the rest of the data as the test set. The training set is split into $K$ subsets of equal size and distributed across $K$ clients in a non-IID fashion. Within each client, $20\%$ of local data are randomly selected as the validation set. Because ChestX-ray14 has a long-tailed label distribution, we choose the 10 most common diseases to ensure that each client can contain labels for all diseases of interest.
6.1.2 Implementation
There are $K=50$ clients in the system and each client is initialized with 64 tokens. Following [51], we use DenseNet121 [52] as the network backbone. We use a standard Adam [46] optimizer with fixed learning rate $10^{-3}$ and batch size 256. We process each CXR with instance normalization [53] and no data augmentation is applied. We use the mean area under the receiver operating characteristic curve over the 10 diseases as both the local validation score and evaluation metric. We set $\epsilon=0.05$ and $\gamma=32$ based on empirical experience. The rest of the implementation details follow Sec. 5.1.2.
6.1.3 Baselines
We consider the same four baselines as in Sec. 5.1.3.
6.2 Results
We run each baseline with 5 random seeds and report both mean and standard deviation under different values of $\eta$. The training results are visualized in Fig. 13.
6.2.1 Performance Comparison
FedAVG w/ block shows competitive performance with FedAVG w/o mal (i.e. missing $\eta=0$) and outperforms FedAVG w/ mal consistently. In addition to the learning curves in Fig. 13, we also report the mean AUROCs for all 4 methods after they fully converge in Tab. 4. When $\eta$ increases, the performance of FedAVG w/ mal drops significantly and becomes more unstable (i.e. missing larger standard deviation). FedAVG w/ block remains robust performance and is only slightly lower than FedAVG w/o mal.
6.2.2 Convergence Analysis
It can be shown that FedAVG w/ block converges faster than both FedAVG w/o mal and FedAVG w/ mal under various values of $\eta$. We hypothesize that the proposed global aggregation mechanism can facilitate federated optimization. Intuitively, this can be explained with gradient descent. FedAVG averages gradients optimized for different directions at different clients, which might not be an optimal global gradient. Under malicious attacks, gradients from malicious clients are intentionally optimized away from the optimal direction, which slows down the training process of FedAVG. However, the proposed consensus mechanism mitigates this issue as it only aggregates when consensus is achieved.
Figure 14: Tokens of honest and malicious clients when $\eta=0.3$ and $\gamma=32$.
6.2.3 Token Analysis
As we use tokens to filter out malicious clients, we plot the average tokens left in honest and malicious clients (e.g. missing Fig. 14). After enough training epochs, honest clients will have more tokens and malicious clients will have fewer tokens. At the end of training, almost all malicious clients do not have enough tokens to stake, i.e. missing they are removed from the FL system. It is worth mentioning that $\gamma$ has only trivial effect on the learning performance but large $\gamma$ can overkill honest clients and small $\gamma$ can cause slow convergence.
6.2.4 Impact of Non-IID Data
Due to the non-IID nature of the medical task, the task setup in this section is more complex than the binary classification task in the previous section. In contrast to Sec. 5, there are two important findings. First, the proposed method is robust under the non-IID setup. Second, surprisingly, while the task is more difficult, the performance gain between the proposed method and the baselines becomes larger.
@@ 7 Conclusion
In this work, we explore an under-explored research direction, namely using FL and blockchain to defend against poisoning attacks. The defense mechanism is twofold. We use on-chain smart contracts to replace the traditional central server and propose a stake-based majority voting mechanism to detect client-side malicious behaviors. We not only provide a solution to the problem of interest, but also show the robustness of the proposed method and provide the first empirical understanding of the problem. Last but not least, the results of this work suggest that the integration of FL and blockchain is an emerging solution to trustworthy ML. We believe that blockchain can not only play an important role in decentralization and the incentivization of participants for real-world FL applications in fields such as finance and medicine, but also can be leveraged to defend against poisoning attacks.
@@ Acknowledgment
The authors would like to thank Shuoying Zhang from FLock.io and Shuhao Zheng from the School of Computer Science, McGill University for the helpful discussion. The authors would also like to thank Yizhe Wen and Xinyang Wang from FLock.io for the experimental simulation.